OpenStudy (anonymous):

solve the system of equations: -0.4x+0.8y = 1.6 0.2x + 0.3y = 2.2

OpenStudy (anonymous):

x=2y−4 x=11−1.5y

OpenStudy (anonymous):

did you rearrange the equation?

OpenStudy (anonymous):

this is what I did for the second one: 0.2x+0.3y=2.2 Since 0.3y does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 0.3y from both sides. 0.2x=−0.3y+2.2 Move all terms not containing x to the right-hand side of the equation. 0.2x=2.2−0.3y Divide each term in the equation by 0.2. 0.2x0.2=2.20.2−0.3y0.2 Simplify the left-hand side of the equation by canceling the common factors. x=2.20.2−0.3y0.2 Simplify the equation.